### v3

# On the lower bound of the curvature exponent on step-two Carnot groups

Abstract.
In this work, we show that there exists a step-two Carnot group on which the new lower bound of the curvature exponent given in [24] can be strictly less than the curvature exponent by studying the convergence of the structure constants of Lie algebra.
Key words and phrases:
Carnot groups, Curvature exponent, Measure contraction property, Sub-Riemannian geometry
2020 Mathematics Subject Classification:
Primary 53C17, 53C23
This project has received funding from (i) the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement GEOSUB, No. 945655); (ii) the JSPS Grant-in-Aid for Early-Career Scientists (No. 24K16928).
1. Introduction
1.1. Measure contraction property
First we recall that measure contraction property is some weak variant of the curvature-dimension condition, which is introduced in [30] or [25]. In the following we assume the underlying space is a length space $(X,d,\mu)$ with negligible cut loci, i.e. a metric measure space with distance function $d$, Borel measure $\mu$, and for every $x\in X$, there exists a negligible set $\mathcal{C}(x)$ and a measurable function $\Phi^{x}:X\setminus\mathcal{C}(x)\times[0,1]\to X$, such that the curve $t\mapsto\Phi^{x}(y,t)$ is the unique length minimizing geodesic joining $x$ and $y$. On such space, we define the set of $s$-intermediate points by
$Z_{s}(x,A)=\{\Phi^{x}(y,s):y\in A\setminus\mathcal{C}(x)\},\quad\forall\,s\in[0,1],x\in X,A\ \mbox{measurable}.$
Definition 1 (Measure contraction property).
A length space $(X,d,\mu)$ with negligible cut loci satisfies $\mathrm{MCP}(K,N)$ for some $K\in\mathbb{R}$ and $N>1$, or $K\leq 0$ and $N=1$, if for every $x\in X$ and measurable $A\subset X$ with $0<\mu(A)<\infty$ (also $A\subset B(x,\pi\sqrt{(N-1)/K})$ if $K>0$), we have
$\mu(Z_{s}(x,A))\geq\int_{A}s\left[\frac{s_{K}(sd(x,y)/\sqrt{N-1})}{s_{K}(d(x,y)/\sqrt{N-1})}\right]^{N-1}d\mu(y).\qquad\forall\,s\in[0,1],$
where we adopt the convention that $0/0=1$ and the term in the square bracket is $1$ for $N=1$. Here and in the following $B(x,r)$ is the open ball centered at $x$ with radius $r$, and the function $s_{K}$ is defined by
$s_{K}(t):=\begin{cases}(1/\sqrt{K})\sin(\sqrt{K}t)\quad&\mbox{if $K>0$},\\ t\quad&\mbox{if $K=0$},\\ (1/\sqrt{-K})\sinh(\sqrt{-K}t)\quad&\mbox{if $K<0$}.\end{cases}$
Here we collect some properties of the measure contraction property. For more details, we refer to [25].
Proposition 2 (Lemma 2.4 and Theorem 4.3 of [25]).
On general length space $(X,d,\mu)$, the following properties hold:
  1. (i)
    $\mathrm{MCP}(K,N)$ implies $\mathrm{MCP}(K^{\prime},N^{\prime})$ for $K^{\prime}\leq K$ and $N^{\prime}\geq N$;
  2. (ii)
    If $\mathrm{MCP}(K,N)$ holds on $(X,d,\mu)$, then $\mathrm{MCP}(a^{-2}K,N)$ holds on $(X,a\,d,b\,\mu)$ for $a,b>0$;
  3. (iii)
    (Generalized Bonnet–Myers theorem) If $\mathrm{MCP}(K,N)$ holds with $K>0$ and $N>1$, then the underlying space is bounded.
1.2. Step-two Carnot groups
Recall that a connected and simply connected Lie group $\mathbb{G}=(\mathbb{G},\cdot)$ is a step-two Carnot group if its left-invariant Lie algebra $\mathfrak{g}$ admits a stratification
$\mathfrak{g}=\mathfrak{g}_{1}\oplus\mathfrak{g}_{2},\quad[\mathfrak{g}_{1},\mathfrak{g}_{1}]=\mathfrak{g}_{2},\quad[\mathfrak{g}_{1},\mathfrak{g}_{2}]=\{0\},$
where $\cdot$$\cdot$] denotes the Lie bracket on $\mathfrak{g}$. Via the Lie group exponential map, we can regard $(\mathbb{G},\cdot)$ as $(\mathfrak{g},\ast)$ with the group operation $\ast$ on $\mathfrak{g}$ given by
(1) $a\ast b:=a+b+\frac{1}{2}[a,b],\qquad\forall\,a,b\in\mathfrak{g}.$
Note that on $(\mathfrak{g},\ast)$ the identity is 0 and the inverse element of $a$ is a-a. For more details, we refer to [9].
Definition 3 (Sub-Riemannian structure).
On a step-two Carnot group $\mathbb{G}$, given an inner product $\langle\cdot,\cdot\rangle$ of $\mathfrak{g}_{1}$, the canonical left-invariant sub-Riemannian structure $(\mathcal{D},g)$ is defined as follows: the horizontal distribution $\mathcal{D}$ (a sub-bundle of the tangent bundle $T\mathbb{G}$ satisfying the bracket generating condition) is generated by $\mathfrak{g}_{1}$ and the metric $g$ on $\mathcal{D}$ is determined by the inner product $\langle\cdot,\cdot\rangle$. To be more precise, let $m:=\dim\mathfrak{g}_{1}$ and assume $\{\mathrm{X}_{1},\ldots,\mathrm{X}_{m}\}$ is an orthonormal basis of $\mathfrak{g}_{1}$ under $\langle\cdot,\cdot\rangle$. Then we have
$\mathcal{D}_{p}:=\mathrm{span}\{\widetilde{\mathrm{X}}_{1}(p),\ldots,\widetilde{\mathrm{X}}_{m}(p)\},$
where
(2) $\widetilde{\mathrm{X}}_{j}(p)=DL_{p}(e)\mathrm{X}_{j},\qquad\forall\,p\in\mathbb{G},$
with $L_{p}(q):=p\cdot q$, $e$ is the identity element on $\mathbb{G}$, and $D$ denotes the differential. Furthermore, $\{\widetilde{\mathrm{X}}_{1}(p),\ldots,\widetilde{\mathrm{X}}_{m}(p)\}$ forms an orthonormal basis of $\mathcal{D}_{p}$ at every point $p\in\mathbb{G}$.
Remark 4.
In this work we regard the Lie algebra $\mathfrak{g}$ as the tangent space of $\mathbb{G}$ at the identity element $e$, namely $T_{e}\mathbb{G}$. So by (2), $\widetilde{\mathrm{X}}_{j}$ are just the left-invariant vector fields on $\mathbb{G}$.
An absolutely continuous path $\gamma:[0,1]\to\mathbb{G}$ is called horizontal if $\dot{\gamma}(s)\in\mathcal{D}_{\gamma(s)}$ for a.e. $s\in[0,1]$, whose length can be calculated by
(3) $\ell(\gamma):=\int_{0}^{1}\sqrt{g(\dot{\gamma}(\tau),\dot{\gamma}(\tau))}\,d\tau.$
Definition 5 (Carnot–Carathéodory distance).
The Carnot–Carathéodory distance between two points $p,q\in\mathbb{G}$ is defined as
$d(p,q):=\inf\{\ell(\gamma):\gamma\mbox{ horizontal},\gamma(0)=p,\gamma(1)=q\}.$
By the celebrated Rashevskii–Chow Theorem (see for example [2, 23]), $d$ is a well-defined finite distance and its induced topology is the same as the manifold topology. In particular, this means that $d$ is continuous with respect to the standard Euclidean topology. Therefore $(\mathbb{G},d)$ is a metric space. A length minimizing geodesic is an horizontal path $\gamma$ such that $\ell(\gamma)=d(\gamma(0),\gamma(1))$.
By definition it is easy to see that the Carnot–Carathéodory distance is left-invariant:
(4) $d(p_{*}\cdot p,p_{*}\cdot q)=d(p,q),\qquad\forall\,p_{*},p,q\in\mathbb{G}.$
For every $\lambda>0$, the dilation $\delta_{\lambda}$ on the Lie algebra $\mathfrak{g}$ is the linear map such that $\delta_{\lambda}(a)=\lambda^{\ell}a$ for $a\in\mathfrak{g}_{\ell}$, $\ell=1,2$. It is actually a Lie algebra automorphism so it induces the dilation on $\mathbb{G}$, which is a Lie group automorphism and we still denote it by $\delta_{\lambda}$ as well. We have the following relation for the Carnot–Carathéodory distance $d$ and the dilation:
(5) $d(\delta_{\lambda}(p),\delta_{\lambda}(q))=\lambda d(p,q),\qquad\forall\,p,q\in\mathbb{G},\lambda>0.$
Finally the Haar measure $\mu$ on $\mathbb{G}$ turns out to be the pushforward measure of the Lebesgue measure $\nu$ on $\mathfrak{g}$ via the Lie group exponential map. With a fixed choice of the Haar measure $\mu$, $(\mathbb{G},d,\mu)$ is a metric measure space. Furthermore, we have
$\mu(\delta_{\lambda}(A))=\lambda^{Q}\mu(A),\qquad\forall\,A\ \mbox{measurable},\lambda>0,$
with homogeneous dimension $Q:=\dim\mathfrak{g}_{1}+2\dim\mathfrak{g}_{2}=m+2\dim\mathfrak{g}_{2}$, which is strictly larger than the topological dimension $n:=\dim\mathfrak{g}_{1}+\dim\mathfrak{g}_{2}=m+\dim\mathfrak{g}_{2}$.
1.3. Measure contraction property on step-two Carnot groups
By [26, Proposition 15] we know that step-two Carnot groups (equipped with the Haar measure) are length spaces with negligible cut loci. From (iii) of Proposition 2, we cannot expect $\mathrm{MCP}(K,N)$ holds on $\mathbb{G}$ with positive $K$ while (ii) of Proposition 2 tells us that the $\mathrm{MCP}(K,N)$ is independent of the choice of the Haar measure $\mu$. Furthermore, from the dilation structure, $(\mathbb{G},d,\mu)$ is isomorphic with $(\mathbb{G},\lambda^{-1}d,\lambda^{-Q}\mu)$. Then (ii) of Proposition 2 also implies that if $\mathrm{MCP}(K,N)$ holds with some negative $K$, then $\mathrm{MCP}(\lambda^{2}\,K,N)$ holds for every $\lambda>0$, which in turn implies $\mathrm{MCP}(0,N)$ holds after taking the limit. However, $\mathrm{MCP}(0,N)$ is stronger than $\mathrm{MCP}(K,N)$ for $K<0$ by (i) of Proposition 2, so we only focus on the study of $\mathrm{MCP}(0,N)$.
Definition 6 (Curvature exponent).
We define $N_{\mathrm{CE}}$, the curvature exponent of $\mathbb{G}$, as the minimum of all $N$ such that $\mathrm{MCP}(0,N)$ holds on $\mathbb{G}$. Equivalently, $N_{\mathrm{CE}}$ is the minimum of all $N$ such that for every $p\in\mathbb{G}$ and measurable $A\subset\mathbb{G}$ with $0<\mu(A)<\infty$, we have
$\mu(Z_{s}(p,A))\geq s^{N}\mu(A),\qquad\forall\,s\in[0,1].$
From the argument above, we know that $\mathrm{MCP}(0,N_{\mathrm{CE}})$ is the best measure contraction property on step-two Carnot group $(\mathbb{G},d,\mu)$ in the sense that if $\mathrm{MCP}(K,N)$ holds, then we must have $K\leq 0$ and $N\geq N_{\mathrm{CE}}$.
The first result about the measure contraction property in the setting of step-two Carnot groups is $N_{\mathrm{CE}}=2Q-n=3n-2m$ on Heisenberg groups by Juillet in [14]. Then similar results are established on corank $1$ Carnot groups and generalized H-type groups in [27] and [6] respectively.
For general step-two Carnot groups, the finiteness of the curvature exponent $N_{\mathrm{CE}}$, is obtained on ideal Carnot groups in [26] and generalized for all step-two Carnot groups:
Theorem 7 (Theorem 1.4 of [4]).
Given a step-two Carnot group $\mathbb{G}$, there exists a constant $N>1$ such that the measure contraction property $\mathrm{MCP}(0,N)$ holds.
Note that it is shown recently in [15, 29, 17] that sub-Riemannian manifolds (equipped with smooth, positive measure) cannot admit any curvature-dimension condition. From Theorem 7, the measure contraction property is a reasonable substitute of such condition in the setting of step-two Carnot groups. Regarding the curvature exponent $N_{\mathrm{CE}}$, in [27], the author proposed the following conjecture.
Conjecture 8.
In the setting of Carnot groups with negligible cut loci, we have $N_{\mathrm{CE}}=N_{\mathrm{GEO}}$.
Here the geodesic dimension $N_{\mathrm{GEO}}$ is defined by
$\sup_{p\in\mathbb{G}}\left[\inf\left\{N>0:\sup_{A\in\mathcal{F}}\limsup_{s\to 0}\frac{\mu(Z_{s}(p,A))}{s^{N}\mu(A)}=\infty\right\}\right],$
where $\mathcal{F}:=\{A\subset\mathbb{G}\text{ bounded, measurable with }0<\mu(A)<\infty\}$. Note that from left-invariance, the outer supremum does not play any role and we can choose any $p$ in the group $\mathbb{G}$, in particular the identity element $e$. For an equivalent definition of the geodesic dimension (at a point), we refer to [1, Definition 5.47]. Similarly, in our setting the point will not play any role. The equivalence is a result of [1, Theorem D] there.
However, Conjecture 8 is not true, which is shown in [24]. More precisely, we found a new lower bound of the curvature exponent, which we denote by $N_{0}$, and a step-two Carnot group on which $N_{\mathrm{CE}}\geq N_{0}>N_{\mathrm{GEO}}$. As a result, there is a gap between $N_{\mathrm{CE}}$ and $N_{\mathrm{GEO}}$. Naturally, we have the following new conjecture.
Conjecture 9.
In the setting of Carnot groups with negligible cut loci, we have $N_{\mathrm{CE}}=N_{0}$.
1.4. Main result
The main target of this work is to show that Conjecture 9 is not true as well. However, it is hard to construct a counterexample directly like the one in [24]. Instead, we illustrate this by showing that the new lower bound $N_{0}$ is not lower semicontinuous w.r.t. the convergence of structure constants of Lie algebra (see Definition 10 and Remark 11 below) while the curvature exponent $N_{\mathrm{CE}}$ is. In the following set $\mathbb{N}=\{1,2,\ldots\}$ and we use $\overline{\mathbb{N}}$ to denote the set $\mathbb{N}\cup\{\infty\}$.
Definition 10.
We say a sequence of Lie algebra $\{(\mathfrak{g},[\cdot,\cdot]_{k})\}_{k\in\mathbb{N}}$ converges to the limit Lie algebra $(\mathfrak{g},[\cdot,\cdot]_{\infty})$ if the followings hold:
1. $\{(\mathfrak{g},[\cdot,\cdot]_{k})\}_{k\in\overline{\mathbb{N}}}$ has the same Lie algebra stratification, i.e.
$\mathfrak{g}=\mathfrak{g}_{1}\oplus\mathfrak{g}_{2},\quad[\mathfrak{g}_{1},\mathfrak{g}_{1}]_{k}=\mathfrak{g}_{2},\quad[\mathfrak{g}_{1},\mathfrak{g}_{2}]_{k}=\{0\},\quad\forall\,k\in\overline{\mathbb{N}};$
2. for any $a,b\in\mathfrak{g}$, we have $a,b]_{k}\to[a,b]_{\infty}$ as $k\to\infty$.
Remark 11.
Fix a basis $\{\mathrm{X}_{1},\ldots,\mathrm{X}_{m}\}$ of $\mathfrak{g}_{1}$ and a basis $\{\mathrm{Y}_{1},\ldots,\mathrm{Y}_{n-m}\}$ of $\mathfrak{g}_{2}$ respectively. For any $1\leq i,j\leq m$, we have
$\mathrm{X}_{i},\mathrm{X}_{j}]_{k}=\sum_{\ell=1}^{n-m}c_{ij,k}^{\ell}\mathrm{Y}_{\ell},\qquad\forall\,k\in\overline{\mathbb{N}}.$
From this equation we know that the convergence of the Lie algebra in Definition 10 is equivalent to the convergence of the structure constants
$c_{ij,k}^{\ell}\to c_{ij,\infty}^{\ell},\qquad\forall\,1\leq i,j\leq m,1\leq\ell\leq n-m$
as $k\to\infty$.
For any fixed $k\in\overline{\mathbb{N}}$, the Lie algebra $\cdot,\cdot]_{k}$ induces a group structure $\ast_{k}$ on $\mathfrak{g}$ by
$a\ast_{k}b:=a+b+\frac{1}{2}[a,b]_{k},\qquad\forall\,a,b\in\mathfrak{g},k\in\overline{\mathbb{N}}.$
In the following we denote $(\mathfrak{g},\ast_{k})$ simply by $\mathbb{G}_{k}$. Furthermore, given some fixed inner product $\langle\cdot,\cdot\rangle$ on $\mathfrak{g}_{1}$ (independent of $k\in\overline{\mathbb{N}}$), we can define a canonical left-invariant sub-Riemannian structure as before and we use $d_{k}$ to denote the Carnot–Carathéodory distance of such sub-Riemannian structure on $\mathbb{G}_{k}$.
For $k\in\overline{\mathbb{N}}$, we define $\mathrm{f}_{k}\in\Gamma(\mathfrak{g}_{1}^{*}\otimes T\mathfrak{g})$ by letting $\mathrm{f}_{k}|_{p}$ to be a linear map from $\mathfrak{g}_{1}$ to $T_{p}\mathfrak{g}$ given by $\mathrm{f}_{k}|_{p}(v):=DL_{p,\mathbb{G}_{k}}(0)v$ for each $p\in\mathfrak{g}$. Here $L_{p,\mathbb{G}_{k}}(q):=p\ast_{k}q$, $k\in\overline{\mathbb{N}}$. Since $\mathfrak{g}$ is a vector space, we can identify $T_{p}\mathfrak{g}$ with $\mathfrak{g}$ itself (like we do for vector spaces, not for Lie groups since we have several group structures here) and after the identification (which is also the bundle trivialization) we have
$\mathrm{f}_{k}|_{p}(v)=v+\frac{1}{2}[p,v]_{k},\qquad\forall\,v\in\mathfrak{g}_{1},p\in\mathfrak{g},k\in\overline{\mathbb{N}}.$
From the expression above it is easy to know that $\{\mathrm{f}_{k}\}_{k\in\overline{\mathbb{N}}}\subset\mathrm{Lip}\Gamma(\mathfrak{g}_{1}^{*}\otimes T\mathfrak{g})$ and $\mathrm{f}_{k}\to\mathrm{f}_{\infty}$ in the topology of $\mathrm{Lip}\Gamma(\mathfrak{g}_{1}^{*}\otimes T\mathfrak{g})$. Here $\mathrm{Lip}\Gamma(\mathfrak{g}_{1}^{*}\otimes T\mathfrak{g})$ denotes the locally Lipschitz sections of $\mathfrak{g}_{1}^{*}\otimes T\mathfrak{g}$. For the precise definition and the topology on such space, we refer to [3, § 2.1]. Now define a continuous varying norm $\mathrm{N}$ on $\mathfrak{g}\times\mathfrak{g}_{1}$ by
$\mathrm{N}(p,v):=\sqrt{\langle v,v\rangle}.$
Then it follows from [3, Corollary 3.19] that $d_{k}=d_{(\mathrm{f}_{k},\mathrm{N})}$ for $k\in\overline{\mathbb{N}}$ and thus the main result there can be applied to our setting. The result is the following.
Proposition 12 ((iv) of Theorem 1.4 of [3]).
Assume $\{(\mathfrak{g},[\cdot,\cdot]_{k})\}_{k\in\mathbb{N}}$ converges to $(\mathfrak{g},[\cdot,\cdot]_{\infty})$ as $k\to\infty$. Then we have $d_{k}(p,q)\to d_{\infty}(p,q)$ uniformly on every compact set of $\mathfrak{g}\times\mathfrak{g}$.
Remark 13.
In the first version of the manuscript Proposition 12 is proved by comparing the endpoint maps of different group structures. For the upper semicontinuity of $k\mapsto d_{k}(p,q)$ w.r.t. the Lie algebra convergence, we used the uniform heat kernel bounds under generalized curvature-dimension inequalities in [7, 8], which is not really needed. As pointed out by the referee, such upper semicontinuity is obtained in [3] with the help of degree theory or quantitative open-mapping type result under suitable smoothness assumption. See also [2, § 3.3.4] for such kind of argument.
One result of Proposition 12 is the pointed measured Gromov–Hausdorff convergence (see [13, Definition 3.24]), which in turn implies the stability of the measure contraction property $\mathrm{MCP}(K,N)$ under the Lie algebra convergence.
Corollary 14 ([13] and Lemma 2.5 of [11]).
Assume $\{(\mathfrak{g},[\cdot,\cdot]_{k})\}_{k\in\mathbb{N}}$ converges to $(\mathfrak{g},[\cdot,\cdot]_{\infty})$ as $k\to\infty$ and fix a Lebesgue measure $\nu$ on $\mathfrak{g}$ (independent of $k\in\overline{\mathbb{N}}$). Then $(\mathbb{G}_{k},d_{k},\nu,0)$ converges to $(\mathbb{G}_{\infty},d_{\infty},\nu,0)$ in the pointed measured Gromov–Hausdorff sense as $k\to\infty$.
As a result, if assume further that for every $k\in\mathbb{N}$, $\mathrm{MCP}(K,N)$ holds on $(\mathbb{G}_{k},d_{k},\nu)$, then $\mathrm{MCP}(K,N)$ also holds on $(\mathbb{G}_{\infty},d_{\infty},\nu)$. In particular, $N_{\mathrm{CE}}$ is lower semicontinuous w.r.t. such Lie algebra convergence.
Now we can disprove Conjecture 9 by showing $N_{0}$ is not lower semicontinuous. In fact, from the construction of our counterexample we can show more. Here is our main result.
Theorem 15.
There exists an ideal Carnot group such that $N_{\mathrm{CE}}>N_{0}$.
Recall that a Carnot group is called ideal if it does not admit non-trivial abnormal minimizing geodesics (see also Definition 21 and Remark 22 below). We also refer to [26, 27] for more details about the definition as well as properties of ideal Carnot groups.
Remark 16.
Recall that in [24] the lower bound $N_{0}$ actually comes from the small time asymptotic behavior of the Jacobian of the sub-Riemannian exponential map. In fact it is also related with the small time asymptotic behavior of sub-Laplacian of the geodesic cost function associated with strongly normal geodesic. For more details we refer to [1, Theorem C], together with [1, Theorem B and Remark 4.11], while a related equivalent characterization of $\mathrm{MCP}(0,N)$ is given in [4, Proposition 2.5]. Thus this corollary says local properties of the Jacobian (or geodesic cost function) cannot characterize the curvature exponent, which is global in nature.
Remark 17.
The phenomenon of the gap between the curvature exponent and the geodesic dimension is also known for the measure contraction property on sub-Finsler Heisenberg groups recently. See the works [12, 10, 11]. Also in this setting, the small time behavior cannot characterize the curvature exponent. See [10, (ii) of Remark 5.13], together with [10, Example 5.14] or [11, Theorem 1.2].
We guess $N_{\mathrm{CE}}$ is actually continuous w.r.t. such Lie algebra convergence and as a result it could be a non integer number while the lower bound $N_{0}$ is always an integer. However, it is hard to prove since we do not know the cut time of every geodesic on step-two Carnot groups. Until now, the cut times are only known on multi-dimensional Heisenberg groups [2, § 13.2], corank $2$ groups [5], free step-two Carnot group with $3$ generators [20], and Reiter–Heisenberg groups [21]. We also refer the interested reader to [28] for the conjecture about the cut locus on free step-two Carnot groups and [22] for the recent pregresses on free step-two Carnot group with $4$ generators.
Conjecture 18.
There exists a step-two Carnot group on which $N_{\mathrm{CE}}$ is not an integer.
1.5. Structure of the paper
In Section 2 we recall the lower bound $N_{0}$ introduced in [24] and the proof of Theorem 15 is provided in Section 3.
2. The lower bound $N_{0}$
Fix a step-two Carnot group $\mathbb{G}$. To give the explicit formula for $N_{0}$ given in [24] in the setting of step-two Carnot groups, we need to first extend the given inner product $\langle\cdot,\cdot\rangle$ on $\mathfrak{g}_{1}$ to the whole Lie algebra $\mathfrak{g}$ in such way that $\mathfrak{g}_{1}$ and $\mathfrak{g}_{2}$ are orthogonal. Although different extension will result in different linear maps and spaces (to be defined below), it turns out that the number $N_{0}$ is the same. Since we only focus on the number $N_{0}$ in this work so we can just fix one such inner product on $\mathfrak{g}$. For the result of this subsection, we refer to [24, § 6.4].
For every $u\in\mathfrak{g}_{2}$, we can define a linear map $J_{u}:\mathfrak{g}_{1}\to\mathfrak{g}_{1}$ by the following formula:
(6) $\langle u,[v,w]\rangle=\langle J_{u}v,w\rangle,\qquad\forall\,v,w\in\mathfrak{g}_{1}.$
From definition $u\mapsto J_{u}$ is linear, continuous, and $J_{u}$ is skew-symmetric since
$\langle J_{u}v,w\rangle=\langle u,[v,w]\rangle=-\langle u,[w,v]\rangle=-\langle J_{u}w,v\rangle=-\langle v,J_{u}w\rangle,\quad\forall\,v,w\in\mathfrak{g}_{1}.$
Furthermore, since $\mathfrak{g}_{1},\mathfrak{g}_{1}]=\mathfrak{g}_{2}$, for $u\in\mathfrak{g}_{2}\setminus\{0\}$ we must have $J_{u}\neq 0$.
Now for fixed $\mathfrak{g}\ni p=\xi+u$ with $\xi\in\mathfrak{g}_{1}$ and $u\in\mathfrak{g}_{2}$, we define
(7) $U^{\ell}(p):=\mathrm{span}\{\xi,J_{u}\xi,\dots,J_{u}^{\ell-1}\xi\}\subset\mathfrak{g}_{1},\quad\forall\,\ell\in\mathbb{N}$
with $U^{0}(p)$ defined to be $\{0\}$, and
(8) $U_{\ell}(p):=\{v\in\mathfrak{g}_{2}:J_{v}(U^{\ell}(p))=\{0\}\}\subset\mathfrak{g}_{2},\quad\forall\,\ell\in\mathbb{N}\cup\{0\}.$
Note that $\{U^{\ell}(p)\}_{\ell\in\mathbb{N}\cup\{0\}}$ is increasing and $\{U_{\ell}(p)\}_{\ell\in\mathbb{N}\cup\{0\}}$ is decreasing. For $\ell\in\mathbb{N}\cup\{0\}$, let $W_{\ell}(p)$ be the orthogonal space of $U_{\ell+1}(p)$ in $U_{\ell}(p)$ and
(9) $W_{\infty}(p):=\left\{v\in\mathfrak{g}_{2}:J_{v}(J_{u}^{\ell}\xi)=0,\forall\,\ell\in\mathbb{N}\cup\{0\}\right\}.$
By definition we have the orthogonal splitting
$\mathfrak{g}_{2}=\bigoplus_{\ell=0}^{\infty}W_{\ell}(p)\oplus W_{\infty}(p).$
Since $\mathfrak{g}_{2}$ is finite dimensional, only finitely many $W_{\ell}(p)$’s are non-trivial. Finally define
(10) $N(p):=\begin{cases}2Q-n+2\sum_{\ell=0}^{\infty}\ell\,\dim(W_{\ell}(p)),\quad&\mbox{if $W_{\infty}(p)=\{0\}$}\\ \infty,\quad&\mbox{if $W_{\infty}(p)\neq\{0\}$}\end{cases}.$
Recall that $Q=\dim\mathfrak{g}_{1}+2\dim\mathfrak{g}_{2}=m+2\dim\mathfrak{g}_{2}$ and $n=\dim\mathfrak{g}_{1}+\dim\mathfrak{g}_{2}=m+\dim\mathfrak{g}_{2}$. The following theorem is the main result of [24] in the setting of step-two Carnot groups.
Remark 19.
The number $N(p)$ defined here is exactly the number $\mathcal{N}_{p}$ in [1, Definition 5.44] after the identification of $\mathfrak{g}=T_{e}\mathbb{G}$ with $\mathfrak{g}^{*}=T_{e}^{*}\mathbb{G}$ induced by the given inner product. For more details, we refer to [24, Remark 1.4 and Remark 6.9].
Theorem 20 (Theorem B and Theorem 6.7 of [24]).
In the setting of step-two Carnot groups, we have
(11) $N_{\mathrm{CE}}\geq N_{0}:=\sup\{N(p):p\in\mathfrak{g}\mbox{ with }N(p)<\infty\}.$
Definition 21 (Métivier group).
We call a step-two Carnot group $\mathbb{G}$ a Métivier group if $J_{v}$ is invertible when $v\neq 0$.
Remark 22.
This kind of groups is first introduced by Métivier in [18]. In the setting of step-two Carnot groups, the condition of Métivier group is equivalent to requiring that there is no non-trivial abnormal length minimizing geodesics (cf. [19, Proposition 3.6] or [16, Equation (3.2)]). As a result, the notion of Métivier groups is equivalent to that of ideal Carnot groups.
By definition, it is easy to prove the following lemma.
Lemma 23.
If $\mathbb{G}$ is a Métivier group, then $N_{0}=2Q-n=\dim\mathfrak{g}_{1}+3\dim\mathfrak{g}_{2}$.
Proof.
From definition we know that if $U^{\ell}(p)\neq\{0\}$, then $U_{\ell}(p)=\{0\}$. For fixed $\mathfrak{g}\ni p=\xi+u$ with $\xi\in\mathfrak{g}_{1}$ and $u\in\mathfrak{g}_{2}$ we know that $U^{\ell}(p)=\{0\}$ if and only if $\ell=0$, or $\xi=0$. As a result, we have
$W_{\infty}(p)=\begin{cases}\mathfrak{g}_{2},\quad&\mbox{if $\xi=0$}\\ \{0\},\quad&\mbox{if $\xi\neq 0$}\end{cases},$
and if $\xi\neq 0$,
$W_{\ell}(p)=\begin{cases}\mathfrak{g}_{2},\quad&\mbox{if $\ell=0$}\\ \{0\},\quad&\mbox{if $\ell\in\mathbb{N}$}\end{cases}.$
From the (10) above, we obtain
$N(p)=\begin{cases}\infty,\quad&\mbox{if $\xi=0$}\\ 2Q-n,\quad&\mbox{if $\xi\neq 0$}\end{cases},$
and thus $N_{0}=2Q-n$ by (11). ∎
3. Proof of Theorem 15
Proof.
On a vector space $\mathfrak{g}=\mathfrak{g}_{1}\oplus\mathfrak{g}_{2}$ with an inner product $\langle\cdot,\cdot\rangle$ w.r.t. which $\mathfrak{g}_{1}$ is orthogonal to $\mathfrak{g}_{2}$, we choose an orthonormal basis of $\mathfrak{g}_{1}$ and $\mathfrak{g}_{2}$ respectively by
$\{\mathrm{X}_{0},\mathrm{X}_{1},\mathrm{X}_{2},\mathrm{X}_{3}\}\quad\mbox{and}\quad\{\mathrm{Y}_{1},\mathrm{Y}_{2},\mathrm{Y}_{3}\}.$
Now we consider the following sequence of Lie algebra structure on $\mathfrak{g}$ defined by the following relations:
$[\mathrm{X}_{0},\mathrm{X}_{1}]_{k}=\mathrm{Y}_{1},\quad$ $[\mathrm{X}_{0},\mathrm{X}_{2}]_{k}=\mathrm{Y}_{2},\quad[\mathrm{X}_{0},\mathrm{X}_{3}]_{k}=\mathrm{Y}_{3},$
$[\mathrm{X}_{1},\mathrm{X}_{2}]_{k}=\frac{1}{k}\mathrm{Y}_{3},\quad$ $[\mathrm{X}_{1},\mathrm{X}_{3}]_{k}=-\frac{1}{k}\mathrm{Y}_{2},\quad[\mathrm{X}_{2},\mathrm{X}_{3}]_{k}=\frac{1}{k}\mathrm{Y}_{1},$
where $k\in\overline{\mathbb{N}}$ and $\frac{1}{\infty}$ is interpreted as 0. In the following for fixed $k\in\overline{\mathbb{N}}$ we use $N_{0,\mathbb{G}_{k}}$, $N_{\mathrm{CE},\mathbb{G}_{k}}$ and $J_{u,\mathbb{G}_{k}}$ to denote the lower bound given in (11), the curvature exponent and the operator defined by (6) on $\mathbb{G}_{k}=(\mathfrak{g},\ast_{k})$ respectively. Note that when $k=\infty$, it is exactly the step-two Carnot group induced by the star graph $K_{1,3}$ in [24, § 7.3] on which $N_{0,\mathbb{G}_{\infty}}=17$. It follows from the construction and Remark 11 that $\{(\mathfrak{g},[\cdot,\cdot]_{k})\}_{k\in\mathbb{N}}$ converges to $(\mathfrak{g},[\cdot,\cdot]_{\infty})$ as $k\to\infty$. Now let $\mathfrak{g}_{2}\ni u=\sum_{\ell=1}^{3}u_{\ell}\mathrm{Y}_{\ell}$, under the basis $\{\mathrm{X}_{0},\mathrm{X}_{1},\mathrm{X}_{2},\mathrm{X}_{3}\}$, we have
$J_{u,\mathbb{G}_{k}}=\begin{pmatrix}0&-u_{1}&-u_{2}&-u_{3}\\ u_{1}&0&-\frac{1}{k}u_{3}&\frac{1}{k}u_{2}\\ u_{2}&\frac{1}{k}u_{3}&0&-\frac{1}{k}u_{1}\\ u_{3}&-\frac{1}{k}u_{2}&\frac{1}{k}u_{1}&0\end{pmatrix}$.
By taking square $J_{u,\mathbb{G}_{k}}^{2}$ we know that $J_{u,\mathbb{G}_{k}}$ is invertible if $u\neq 0$ and $k\in\mathbb{N}$, which means $\mathbb{G}_{k}(k\in\mathbb{N})$ is Métivier, or equivalently ideal by Remark 22. As a result, we know that $N_{0,\mathbb{G}_{k}}=13$ for $k\in\mathbb{N}$ by Lemma 23. Assume on the contrary that $N_{\mathrm{CE}}=N_{0}$ holds for all ideal Carnot groups. Then we know $N_{\mathrm{CE},\mathbb{G}_{k}}=N_{0,\mathbb{G}_{k}}=13$ for $k\in\mathbb{N}$. It follows from Proposition 12 that
$17=N_{0,\mathbb{G}_{\infty}}\leq N_{\mathrm{CE},\mathbb{G}_{\infty}}\leq\liminf_{k\to\infty}N_{\mathrm{CE},\mathbb{G}_{k}}=\liminf_{k\to\infty}N_{0,\mathbb{G}_{k}}=13,$
which leads to a contradiction. We complete the proof. In particular, we have showed that $N_{0}$ is not lower semicontinuous w.r.t. the Lie algebra convergence.
Acknowledgements
The author would like to thank the anonymous referee for many valuable suggestions and remarks, especially the reference to the paper [3], which reduces the length of the proof significantly. The author would also like to thank Kenshiro Tashiro for fruitful discussions.